How many moles of water would be produced by the complete combustion of one mole of benzene, C6H6, in excess oxygen?
1
Expert's answer
2020-04-03T11:52:15-0400

2C6H6 + 15O2 = 12CO2 + 6H2O

n(H2O) = (1*6)/2 = 3 moles